vertex standard form?
1. y=x^2-4x-5 ?put this function in the form y=a(x-h)+k.

2. What is the line of symmetry?

3. Graph th function using the equation in part a, explain why it is not necessary to plot points to graph when using y=a(x-h)^+k? Explanation of graphing?
4. In a few sentances describe how this graph compares to the graph of y=x^?

Answer by mathnuts(1)

1. y = x^2 - 4x - 5
= x^2 - 4x + 4 - 4 - 5
= (x^2 - 4x + 4) -4 - 5
y = (x-2)^2 - 9 Here, the vertex is at (2, -9)
2. x = 2 (the vertical line that goes through the vertex)
3. The graph of this function is called a parabola. To graph a parabola, use
its vertex (in this case, (2, -9) ) and y-intercept (y = -5). The y-intercept is the constant in the given equation. Use the fact that the vertical line through the vertex is a line of symmetry. Plot the point corresponding to the y-intercept that lies on the other side of the vertical line x = 2.

4. The graph is translated 2 units to the right and 9 units down.